# Bootstrapping the 3d Ising Stress Tensor

### v1

@@ 6 Conclusions
One amusing interpretation of our determinations of $\langle TTT\rangle$ OPE coefficients is that the 3d Ising CFT is “$n_{F}/c_{T}=1.3833683(35)\%$ fermionic”. We may wonder how this percentage compares against other interacting CFTs, such as the critical $O(N)$ vector models or CFTs built from interacting fermions or gauge fields. In particular, how these OPE coefficients may vary across different families of CFTs could well be nontrivial in light of conjectured bosonization/fermionization dualities Seiberg:2016gmd ; see Turner:2019wnh for an excellent pedagogical review. Bootstrapping other mixed systems involving the stress tensor therefore would be an interesting line of inquiry. We leave these exciting directions to future work.
@@ Appendix A Three-point structures
In this appendix, we record our choice of three-point structures for correlators of $\sigma,\epsilon,T$ in the 3d Ising CFT. We follow the conventions of Erramilli:2022kgp , where we provide a three-point structure for each ordered triple of operators that can appear in a three-point correlator $\langle\mathcal{O}_{i}\mathcal{O}_{j}\mathcal{O}_{k}\rangle$. These structures must be chosen so that the OPE coefficients $\lambda^{a}_{ijk}$ are functions of the unordered triple $(i,j,k)$. [^25] In practice this means that the structures are related by permutations. By providing the functions explicitly, we allow the code that prepares conformal block expansions to be independent of the details of how permutations act on local operators. When all three operators are external (i.e. $\sigma,\epsilon,$ or $T$), we specify structures for all unique ordered triples. When one of the operators is internal, we choose conventions where the internal operator appears last, so we only specify structures for up to two unique orderings.
@@
$\begin{array}[]{|l|l|l|}\hline\cr\textrm{OPE coefficient}&\textrm{ordered triple $\langle\mathcal{O}_{i}\mathcal{O}_{j}\mathcal{O}_{j}\rangle$}&\textrm{three-point structure}\\ \hline\cr\lambda_{B}&\langle TTT\rangle&\texttt{StressTensorB}\\ &\langle T\sigma\sigma\rangle&-\Delta_{\sigma}|2,2\rangle\\ &\langle\sigma T\sigma\rangle&-\Delta_{\sigma}|2,2\rangle\\ &\langle\sigma\sigma T\rangle&-\Delta_{\sigma}|0,2\rangle\\ &\langle T\epsilon\epsilon\rangle&-\Delta_{\epsilon}|2,2\rangle\\ &\langle\epsilon T\epsilon\rangle&-\Delta_{\epsilon}|2,2\rangle\\ &\langle\epsilon\epsilon T\rangle&-\Delta_{\epsilon}|0,2\rangle\\ \hline\cr\lambda_{F}&\langle TTT\rangle&\texttt{StressTensorF}\\ &\langle T\sigma\sigma\rangle&-\Delta_{\sigma}|2,2\rangle\\ &\langle\sigma T\sigma\rangle&-\Delta_{\sigma}|2,2\rangle\\ &\langle\sigma\sigma T\rangle&-\Delta_{\sigma}|0,2\rangle\\ &\langle T\epsilon\epsilon\rangle&-\Delta_{\epsilon}|2,2\rangle\\ &\langle\epsilon T\epsilon\rangle&-\Delta_{\epsilon}|2,2\rangle\\ &\langle\epsilon\epsilon T\rangle&-\Delta_{\epsilon}|0,2\rangle\\ \hline\cr\lambda_{\sigma\sigma\epsilon}&\langle\sigma\sigma\epsilon\rangle&|0,0\rangle\\ &\langle\sigma\epsilon\sigma\rangle&|0,0\rangle\\ &\langle\epsilon\sigma\sigma\rangle&|0,0\rangle\\ \hline\cr\lambda_{\epsilon\epsilon\epsilon}&\langle\epsilon\epsilon\epsilon\rangle&|0,0\rangle\\ \hline\cr\lambda_{TT\epsilon}&\langle TT\epsilon\rangle&-\frac{16\Delta_{\epsilon}}{2+\Delta_{\epsilon}}\ \texttt{ScalarParityEven}\\ &\langle T\epsilon T\rangle&\frac{6}{\Delta_{\epsilon}(2+\Delta_{\epsilon})}\ \texttt{TSOParityEven}_{2}\\ &\langle\epsilon TT\rangle&\frac{6}{\Delta_{\epsilon}(2+\Delta_{\epsilon})}\ \texttt{STOParityEven}_{2}\\ \hline\cr\end{array}$ (85)
@@
$\begin{array}[]{|l|l|}\hline\cr\langle TT\mathcal{O}\rangle&\texttt{ScalarParityEven}\\ \hline\cr\langle\sigma\sigma\mathcal{O}\rangle&|0,0\rangle\\ \hline\cr\langle\epsilon\epsilon\mathcal{O}\rangle&|0,0\rangle\\ \hline\cr\end{array}$ (89)
@@
$\begin{array}[]{|l|l|}\hline\cr\langle TT\mathcal{O}\rangle&\texttt{Spin2ParityEven}\\ \hline\cr\langle\sigma\sigma\mathcal{O}\rangle&|0,2\rangle\\ \hline\cr\langle\epsilon\epsilon\mathcal{O}\rangle&|0,2\rangle\\ \hline\cr\langle T\epsilon\mathcal{O}\rangle&\texttt{TSOParityEven}_{2}\\ \langle\epsilon T\mathcal{O}\rangle&\texttt{STOParityEven}_{2}\\ \hline\cr\end{array}$ (95)
@@
$\begin{array}[]{|l|l|}\hline\cr\langle T\epsilon\mathcal{O}\rangle&\texttt{TSOParityEven}_{j}\\ \langle\epsilon T\mathcal{O}\rangle&\texttt{STOParityEven}_{j}\\ \hline\cr\end{array}$ (98)
@@
$\begin{array}[]{|l|l|}\hline\cr\langle TT\mathcal{O}\rangle&\texttt{GenericParityEven1}_{j}\\ \hline\cr\langle TT\mathcal{O}\rangle&\texttt{GenericParityEven2}_{j}\\ \hline\cr\langle\sigma\sigma\mathcal{O}\rangle&|0,j\rangle\\ \hline\cr\langle\epsilon\epsilon\mathcal{O}\rangle&|0,j\rangle\\ \hline\cr\langle T\epsilon\mathcal{O}\rangle&\texttt{TSOParityEven}_{j}\\ \langle\epsilon T\mathcal{O}\rangle&\texttt{STOParityEven}_{j}\\ \hline\cr\end{array}$ (105)
@@
$\begin{array}[]{|l|l|}\hline\cr\langle TT\mathcal{O}\rangle&\texttt{ScalarParityOdd}\\ \hline\cr\end{array}$ (107)
@@
$\begin{array}[]{|l|l|}\hline\cr\langle TT\mathcal{O}\rangle&\texttt{Spin2ParityOdd}\\ \hline\cr\langle T\epsilon\mathcal{O}\rangle&\texttt{TSOParityOdd}_{2}\\ \langle\epsilon T\mathcal{O}\rangle&\texttt{STOParityOdd}_{2}\\ \hline\cr\end{array}$ (111)
@@
$\begin{array}[]{|l|l|}\hline\cr\langle T\epsilon\mathcal{O}\rangle&\texttt{TSOParityOdd}_{3}\\ \langle\epsilon T\mathcal{O}\rangle&\texttt{STOParityOdd}_{3}\\ \hline\cr\end{array}$ (114)
@@
$\begin{array}[]{|l|l|}\hline\cr\langle TT\mathcal{O}\rangle&\texttt{GenericParityOdd}_{j}\\ \hline\cr\langle T\epsilon\mathcal{O}\rangle&\texttt{TSOParityOdd}_{j}\\ \langle\epsilon T\mathcal{O}\rangle&\texttt{STOParityOdd}_{j}\\ \hline\cr\end{array}$ (118)
@@
$\begin{array}[]{|l|l|}\hline\cr\langle\sigma\epsilon\mathcal{O}\rangle&|0,0\rangle\\ \langle\epsilon\sigma\mathcal{O}\rangle&|0,0\rangle\\ \hline\cr\end{array}$ (121)
@@
$\begin{array}[]{|l|l|}\hline\cr\langle\sigma\epsilon\mathcal{O}\rangle&|0,1\rangle\\ \langle\epsilon\sigma\mathcal{O}\rangle&-|0,1\rangle\\ \hline\cr\end{array}$ (124)
@@
$\begin{array}[]{|l|l|}\hline\cr\langle T\sigma\mathcal{O}\rangle&\texttt{TSOParityEven}_{j}\\ \langle\sigma T\mathcal{O}\rangle&\texttt{STOParityEven}_{j}\\ \hline\cr\langle\sigma\epsilon\mathcal{O}\rangle&|0,j\rangle\\ \langle\epsilon\sigma\mathcal{O}\rangle&(-1)^{j}|0,j\rangle\\ \hline\cr\end{array}$ (129)
@@
$\begin{array}[]{|l|l|}\hline\cr\langle T\sigma\mathcal{O}\rangle&\texttt{TSOParityOdd}_{j}\\ \langle\sigma T\mathcal{O}\rangle&\texttt{STOParityOdd}_{j}\\ \hline\cr\end{array}$ (132)
@@ Appendix B Crossing equations for the $T\sigma\epsilon$ system
[…]
@@
[…]
@@
00 0 $=\left.\left(g_{TTTT}^{\langle 0,0,2,2\rangle^{+}}-g_{TTTT}^{\langle 2,0,0,2\rangle^{+}}\right)\right|_{*},$
00 0 $=\left.\left(g_{TTTT}^{\langle 0,-1,1,2\rangle^{+}}-g_{TTTT}^{\langle 1,-1,0,2\rangle^{+}}\right)\right|_{*},$
00 0 $=\left.\left(g_{TTTT}^{\langle-1,-1,1,1\rangle^{+}}-g_{TTTT}^{\langle 1,-1,-1,1\rangle^{+}}\right)\right|_{*},$
00 0 $=\left.\left(g_{TTTT}^{\langle-1,1,0,2\rangle^{+}}-g_{TTTT}^{\langle 0,1,-1,2\rangle^{+}}\right)\right|_{*}.$ (137)
@@
[…]
@@
[…]
@@
00 0 $=\left.\left(g_{TTT\epsilon}^{\langle-2,2,2,0\rangle^{+}}-g_{TTT\epsilon}^{\langle 2,2,-2,0\rangle^{+}}\right)\right|_{*}.$ (140)
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]

### v2

@@ 6 Conclusions
One amusing interpretation of our determinations of $\langle TTT\rangle$ OPE coefficients is that the 3d Ising CFT is $n_{F}/(n_{F}+n_{B})=1.3833683(35)\%$ fermionic”. We may wonder how this percentage compares against other interacting CFTs, such as the critical $O(N)$ vector models or CFTs built from interacting fermions or gauge fields. In particular, how these OPE coefficients may vary across different families of CFTs could well be nontrivial in light of conjectured bosonization/fermionization dualities Seiberg:2016gmd ; see Turner:2019wnh for an excellent pedagogical review. Bootstrapping other mixed systems involving the stress tensor therefore would be an interesting line of inquiry. We leave these exciting directions to future work.
@@ Appendix A Three-point structures
In this appendix, we record our choice of three-point structures for correlators of $\sigma,\epsilon,T$ in the 3d Ising CFT. We follow the conventions of Erramilli:2022kgp , where we provide a three-point structure for each ordered triple of operators that can appear in a three-point correlator $\langle\mathcal{O}_{i}\mathcal{O}_{j}\mathcal{O}_{k}\rangle$. These structures must be chosen so that the OPE coefficients $\lambda^{a}_{ijk}$ are functions of the unordered triple (i,j,k)(i,j,k). [^25] In practice this means that the structures are related by permutations. By providing the functions explicitly, we allow the code that prepares conformal block expansions to be independent of the details of how permutations act on local operators. When all three operators are external (i.e. $\sigma,\epsilon,$ or $T$), we specify structures for all unique ordered triples. When one of the operators is internal, we choose conventions where the internal operator appears last, so we only specify structures for up to two unique orderings.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ Appendix B Crossing equations for the $T\sigma\epsilon$ system
$0$ $=\left.\partial_{w}^{m}\partial_{s}^{n}\left(g_{TTTT}^{\langle 2,2,2,2\rangle^{+}}\right)\right|_{*},$ $(\textrm{$m$ odd})$
$0$ $=\left.\partial_{w}^{m}\partial_{s}^{n}\left(g_{TTTT}^{\langle 1,1,1,1\rangle^{+}}\right)\right|_{*},$ $(\textrm{$m$ odd})$
$0$ $=\left.\partial_{w}^{m}\partial_{s}^{n}\left(g_{TTTT}^{\langle 1,1,2,2\rangle^{+}}+g_{TTTT}^{\langle 2,1,1,2\rangle^{+}}\right)\right|_{*},$ $(\textrm{$m$ odd})$
$0$ $=\left.\partial_{w}^{m}\partial_{s}^{n}\left(g_{TTTT}^{\langle 1,1,2,2\rangle^{+}}-g_{TTTT}^{\langle 2,1,1,2\rangle^{+}}\right)\right|_{*},$ $(\textrm{$m$ even})$
$0$ $=\left.\partial_{w}^{m}\partial_{s}^{n}\left(g_{TTTT}^{\langle 1,2,1,2\rangle^{+}}\right)\right|_{*}.$ $(\textrm{$m$ odd})$ (135)
@@
$0$ $=\left.\partial_{w}^{m}\left(g_{TTTT}^{\langle 0,0,0,0\rangle^{+}}\right)\right|_{*},$ $(\textrm{$m$ odd})$
$0$ $=\left.\partial_{w}^{m}\left(g_{TTTT}^{\langle 0,1,0,1\rangle^{+}}\right)\right|_{*},$ $(\textrm{$m$ odd})$
$0$ $=\left.\partial_{w}^{m}\left(g_{TTTT}^{\langle 0,2,0,2\rangle^{+}}\right)\right|_{*},$ $(\textrm{$m$ odd})$
$0$ $=\left.\partial_{w}^{m}\left(g_{TTTT}^{\langle 0,1,1,2\rangle^{+}}+g_{TTTT}^{\langle 1,1,0,2\rangle^{+}}\right)\right|_{*},$ $(\textrm{$m$ odd})$
$0$ $=\left.\partial_{w}^{m}\left(g_{TTTT}^{\langle 0,1,1,2\rangle^{+}}-g_{TTTT}^{\langle 1,1,0,2\rangle^{+}}\right)\right|_{*},$ $(\textrm{$m$ even})$
$0$ $=\left.\partial_{w}^{m}\left(g_{TTTT}^{\langle 0,0,1,1\rangle^{+}}+g_{TTTT}^{\langle 1,0,0,1\rangle^{+}}\right)\right|_{*},$ $(\textrm{$m$ odd})$
$0$ $=\left.\partial_{w}^{m}\left(g_{TTTT}^{\langle 0,0,1,1\rangle^{+}}-g_{TTTT}^{\langle 1,0,0,1\rangle^{+}}\right)\right|_{*},$ $(\textrm{$m$ even})$
$0$ $=\left.\partial_{w}^{m}\left(g_{TTTT}^{\langle-1,0,0,1\rangle^{+}}+g_{TTTT}^{\langle 0,0,-1,1\rangle^{+}}\right)\right|_{*},$ $(\textrm{$m$ odd})$
$0$ $=\left.\partial_{w}^{m}\left(g_{TTTT}^{\langle-1,0,0,1\rangle^{+}}-g_{TTTT}^{\langle 0,0,-1,1\rangle^{+}}\right)\right|_{*}.$ $(\textrm{$m$ even})$ (136)
@@
$0$ $=\left.\left(g_{TTTT}^{\langle 0,0,2,2\rangle^{+}}-g_{TTTT}^{\langle 2,0,0,2\rangle^{+}}\right)\right|_{*},$
$0$ $=\left.\left(g_{TTTT}^{\langle 0,-1,1,2\rangle^{+}}-g_{TTTT}^{\langle 1,-1,0,2\rangle^{+}}\right)\right|_{*},$
$0$ $=\left.\left(g_{TTTT}^{\langle-1,-1,1,1\rangle^{+}}-g_{TTTT}^{\langle 1,-1,-1,1\rangle^{+}}\right)\right|_{*},$
$0$ $=\left.\left(g_{TTTT}^{\langle-1,1,0,2\rangle^{+}}-g_{TTTT}^{\langle 0,1,-1,2\rangle^{+}}\right)\right|_{*}.$ (137)
@@
$0$ $=\left.\partial_{w}^{m}\partial_{s}^{n}\left(g_{TTT\epsilon}^{\langle 2,2,2,0\rangle^{+}}\right)\right|_{*},$ $(\textrm{$m$ odd})$
$0$ $=\left.\partial_{w}^{m}\partial_{s}^{n}\left(g_{TTT\epsilon}^{\langle 1,2,1,0\rangle^{+}}\right)\right|_{*},$ $(\textrm{$m$ odd})$
$0$ $=\left.\partial_{w}^{m}\partial_{s}^{n}\left(g_{TTT\epsilon}^{\langle 1,1,2,0\rangle^{+}}+g_{TTT\epsilon}^{\langle 2,1,1,0\rangle^{+}}\right)\right|_{*},$ $(\textrm{$m$ odd})$
$0$ $=\left.\partial_{w}^{m}\partial_{s}^{n}\left(g_{TTT\epsilon}^{\langle 1,1,2,0\rangle^{+}}-g_{TTT\epsilon}^{\langle 2,1,1,0\rangle^{+}}\right)\right|_{*}.$ $(\textrm{$m$ even})$ (138)
@@
$0$ $=\left.\partial_{w}^{m}\left(g_{TTT\epsilon}^{\langle 0,0,2,0\rangle^{+}}+g_{TTT\epsilon}^{\langle 2,0,0,0\rangle^{+}}\right)\right|_{*},$ $(\textrm{$m$ odd})$
$0$ $=\left.\partial_{w}^{m}\left(g_{TTT\epsilon}^{\langle 0,0,2,0\rangle^{+}}-g_{TTT\epsilon}^{\langle 2,0,0,0\rangle^{+}}\right)\right|_{*},$ $(\textrm{$m$ even})$
$0$ $=\left.\partial_{w}^{m}\left(g_{TTT\epsilon}^{\langle 0,2,0,0\rangle^{+}}\right)\right|_{*}.$ $(\textrm{$m$ odd})$ (139)
@@
$0$ $=\left.\left(g_{TTT\epsilon}^{\langle-2,2,2,0\rangle^{+}}-g_{TTT\epsilon}^{\langle 2,2,-2,0\rangle^{+}}\right)\right|_{*}.$ (140)
@@
$0$ $=\left.\partial_{w}^{m}\partial_{s}^{n}\left(g_{TT\sigma\sigma}^{\langle 2,2,0,0\rangle^{+}}+g_{\sigma TT\sigma}^{\langle 0,2,2,0\rangle^{+}}\right)\right|_{*},$ $(\textrm{$m$ odd})$
$0$ $=\left.\partial_{w}^{m}\partial_{s}^{n}\left(g_{TT\sigma\sigma}^{\langle 2,2,0,0\rangle^{+}}-g_{\sigma TT\sigma}^{\langle 0,2,2,0\rangle^{+}}\right)\right|_{*},$ $(\textrm{$m$ even})$
$0$ $=\left.\partial_{w}^{m}\partial_{s}^{n}\left(g_{TT\sigma\sigma}^{\langle 1,1,0,0\rangle^{+}}+g_{\sigma TT\sigma}^{\langle 0,1,1,0\rangle^{+}}\right)\right|_{*},$ $(\textrm{$m$ odd})$
$0$ $=\left.\partial_{w}^{m}\partial_{s}^{n}\left(g_{TT\sigma\sigma}^{\langle 1,1,0,0\rangle^{+}}-g_{\sigma TT\sigma}^{\langle 0,1,1,0\rangle^{+}}\right)\right|_{*},$ $(\textrm{$m$ even})$
$0$ $=\left.\partial_{w}^{m}\partial_{s}^{n}\left(g_{\sigma T\sigma T}^{\langle 0,2,0,2\rangle^{+}}\right)\right|_{*},$ $(\textrm{$m$ odd})$
$0$ $=\left.\partial_{w}^{m}\partial_{s}^{n}\left(g_{\sigma T\sigma T}^{\langle 0,1,0,1\rangle^{+}}\right)\right|_{*}.$ $(\textrm{$m$ odd})$ (141)
@@
$0$ $=\left.\partial_{w}^{m}\left(g_{TT\sigma\sigma}^{\langle 0,0,0,0\rangle^{+}}+g_{\sigma TT\sigma}^{\langle 0,0,0,0\rangle^{+}}\right)\right|_{*},$ $(\textrm{$m$ odd})$
$0$ $=\left.\partial_{w}^{m}\left(g_{TT\sigma\sigma}^{\langle 0,0,0,0\rangle^{+}}-g_{\sigma TT\sigma}^{\langle 0,0,0,0\rangle^{+}}\right)\right|_{*},$ $(\textrm{$m$ even})$
$0$ $=\left.\partial_{w}^{m}\left(g_{\sigma T\sigma T}^{\langle 0,0,0,0\rangle^{+}}\right)\right|_{*}.$ $(\textrm{$m$ odd})$ (142)
@@
$0$ $=\left.\partial_{w}^{m}\partial_{s}^{n}\left(g_{TT\epsilon\epsilon}^{\langle 2,2,0,0\rangle^{+}}+g_{\epsilon TT\epsilon}^{\langle 0,2,2,0\rangle^{+}}\right)\right|_{*},$ $(\textrm{$m$ odd})$
$0$ $=\left.\partial_{w}^{m}\partial_{s}^{n}\left(g_{TT\epsilon\epsilon}^{\langle 2,2,0,0\rangle^{+}}-g_{\epsilon TT\epsilon}^{\langle 0,2,2,0\rangle^{+}}\right)\right|_{*},$ $(\textrm{$m$ even})$
$0$ $=\left.\partial_{w}^{m}\partial_{s}^{n}\left(g_{TT\epsilon\epsilon}^{\langle 1,1,0,0\rangle^{+}}+g_{\epsilon TT\epsilon}^{\langle 0,1,1,0\rangle^{+}}\right)\right|_{*},$ $(\textrm{$m$ odd})$
$0$ $=\left.\partial_{w}^{m}\partial_{s}^{n}\left(g_{TT\epsilon\epsilon}^{\langle 1,1,0,0\rangle^{+}}-g_{\epsilon TT\epsilon}^{\langle 0,1,1,0\rangle^{+}}\right)\right|_{*},$ $(\textrm{$m$ even})$
$0$ $=\left.\partial_{w}^{m}\partial_{s}^{n}\left(g_{\epsilon T\epsilon T}^{\langle 0,2,0,2\rangle^{+}}\right)\right|_{*},$ $(\textrm{$m$ odd})$
$0$ $=\left.\partial_{w}^{m}\partial_{s}^{n}\left(g_{\epsilon T\epsilon T}^{\langle 0,1,0,1\rangle^{+}}\right)\right|_{*}.$ $(\textrm{$m$ odd})$ (143)
@@
$0$ $=\left.\partial_{w}^{m}\left(g_{TT\epsilon\epsilon}^{\langle 0,0,0,0\rangle^{+}}+g_{\epsilon TT\epsilon}^{\langle 0,0,0,0\rangle^{+}}\right)\right|_{*},$ $(\textrm{$m$ odd})$
$0$ $=\left.\partial_{w}^{m}\left(g_{TT\epsilon\epsilon}^{\langle 0,0,0,0\rangle^{+}}-g_{\epsilon TT\epsilon}^{\langle 0,0,0,0\rangle^{+}}\right)\right|_{*},$ $(\textrm{$m$ even})$
$0$ $=\left.\partial_{w}^{m}\left(g_{\epsilon T\epsilon T}^{\langle 0,0,0,0\rangle^{+}}\right)\right|_{*}.$ $(\textrm{$m$ odd})$ (144)
@@
$0$ $=\left.\partial_{w}^{m}\partial_{s}^{n}\left(g_{\epsilon T\epsilon\epsilon}^{\langle 0,2,0,0\rangle^{+}}\right)\right|_{*}.$ $(\textrm{$m$ odd})$ (145)
@@
$0$ $=\left.\partial_{w}^{m}\partial_{s}^{n}\left(g_{\sigma T\epsilon\sigma}^{\langle 0,2,0,0\rangle^{+}}+g_{\epsilon T\sigma\sigma}^{\langle 0,2,0,0\rangle^{+}}\right)\right|_{*},$ $(\textrm{$m$ odd})$
$0$ $=\left.\partial_{w}^{m}\partial_{s}^{n}\left(g_{\sigma T\epsilon\sigma}^{\langle 0,2,0,0\rangle^{+}}-g_{\epsilon T\sigma\sigma}^{\langle 0,2,0,0\rangle^{+}}\right)\right|_{*},$ $(\textrm{$m$ even})$
$0$ $=\left.\partial_{w}^{m}\partial_{s}^{n}\left(g_{\sigma T\sigma\epsilon}^{\langle 0,2,0,0\rangle^{+}}\right)\right|_{*}.$ $(\textrm{$m$ odd})$ (146)
@@
$0$ $=\left.\partial_{w}^{m}\partial_{s}^{n}\left(g_{\sigma\sigma\sigma\sigma}^{\langle 0,0,0,0\rangle^{+}}\right)\right|_{*},$ $(\textrm{$m$ odd})$
$0$ $=\left.\partial_{w}^{m}\partial_{s}^{n}\left(g_{\epsilon\epsilon\epsilon\epsilon}^{\langle 0,0,0,0\rangle^{+}}\right)\right|_{*},$ $(\textrm{$m$ odd})$
$0$ $=\left.\partial_{w}^{m}\partial_{s}^{n}\left(g_{\sigma\sigma\epsilon\epsilon}^{\langle 0,0,0,0\rangle^{+}}+g_{\epsilon\sigma\sigma\epsilon}^{\langle 0,0,0,0\rangle^{+}}\right)\right|_{*},$ $(\textrm{$m$ odd})$
$0$ $=\left.\partial_{w}^{m}\partial_{s}^{n}\left(g_{\sigma\sigma\epsilon\epsilon}^{\langle 0,0,0,0\rangle^{+}}-g_{\epsilon\sigma\sigma\epsilon}^{\langle 0,0,0,0\rangle^{+}}\right)\right|_{*},$ $(\textrm{$m$ even})$
$0$ $=\left.\partial_{w}^{m}\partial_{s}^{n}\left(g_{\sigma\epsilon\sigma\epsilon}^{\langle 0,0,0,0\rangle^{+}}\right)\right|_{*}.$ $(\textrm{$m$ odd})$ (147)